# arXiv:2603.02804v1[quant-ph] 03 Mar 2026

### 03 Mar 2026 v1

@@ 4 Numerical experiments
In this section, to evaluate the performance of our proposed method, implemented with Triton [triton2019] and PyTorch [Ansel_PyTorch_2_Faster_2024], we measured the average execution time over $1,000$ runs using the benchmarking utility triton.testing.do_bench in Triton [triton2019]. We also measured peak memory usage using torch.cuda.max_memory_allocated from a single measurement, since peak memory usage remains consistent across multiple executions. Note that although our implementation supports both single precision (float32/complex64) and double precision (float64/complex128), we use float32 (complex64) in our numerical experiments to focus on performance evaluation for practical QML tasks, which often have severe memory constraints. Furthermore, throughout our numerical experiments, we apply quantum circuits to randomly initialized quantum states $|\psi_{0}\rangle$ rather than $|0\rangle^{\otimes n}$ in order to avoid artificial speedups caused by sparsity and to reflect the actual computational load and memory bandwidth utilization.
@@ 4.1 The effect of gate fusion
In this subsection, we investigate the effect of the gate fusion on a quantum circuit consisting of consecutive single-qubit gates, as shown in Fig. 3. Using a $20$-qubit system and a batch size of $100$, we focus on how memory usage and execution time depend on the number of fused gates in the forward and backward paths.
First, we measured the mean execution time per fused gate for a quantum circuit where $m$ consecutive single-qubit gates (Rx, Ry, and Rz) are applied to each qubit during the forward path. As shown in Fig. 4(a), we can see that our proposed method allows us to drastically accelerate the classical simulation, compared with the PyTorch native implementation, which applies each gate sequentially. In our numerical experiments, since we apply the gates in the order of Rx, Ry, and Rz, the execution time of the PyTorch native implementation gradually increases, especially when applying the diagonal matrix Rz gate. Furthermore, Fig. 4(a) shows that the execution time of our proposed method increases almost linearly with $m$, but much more gradually than that of the PyTorch native implementation.
As shown in Fig. 4(b), the peak memory usage is almost the same regardless of the number of gates to be fused. In the forward path, the peak memory usage of our proposed method (Triton fused) is the lowest among the three implementations. The total peak memory usage is $2,500$ MB, roughly corresponding to three batched state vectors ($3\times 800$ MB). It appears that the input, output, and temporary batched state vectors are held, and the remaining $100$ MB is internal overhead. Similarly, the Triton fused (mem_save) implementation used an additional $400$ MB, caused by storing the batched state vector converted to bfloat16 for the backward path. The PyTorch native implementation consumes approximately $800$ MB more memory than the Triton fused implementation, corresponding to a single batched state vector. This suggests that the PyTorch native implementation temporarily keeps a copy of the batched state vectors during the forward path, likely due to explicitly cloning them for safety. Since PyTorch’s automatic differentiation requires out-of-place operations on inputs to preserve the computation graph, our implementation is considered to be memory-efficient.
@@
(a) Execution time per fused gate
@@
Figure 4: (a) Execution time per fused gate and (b) Total peak memory usage when applying $m$ consecutive single-qubit gates on each qubit during the forward path
Next, to evaluate the performance of the backward function itself, we apply the same quantum circuit as above, shown in Fig. 3, and measure the mean execution time per fused kernel. From the results shown in Fig. 5(a), for consumer GPUs such as the RTX 5070 and RTX 4090, the computation time of applying a few gates to the input quantum state is almost constant, i.e., the recomputation time is negligible because the memory access latency to read quantum states and adjoint states from global memory is dominant. As the number of gates to be fused further increases, the computation time begins to scale linearly. On the other hand, for data center GPUs with large memory bandwidth such as the GH200, the computation time increases linearly with the number of gates to be fused from the beginning. As shown in Fig. 5(b), the peak memory usage linearly increases with the number of gates to be fused, and only $30$ KB is increased (from $2516.585$ MB to $2516.613$ MB). This indicates that the computation of the fused gates in the backward path is processed efficiently while mitigating the use of temporary variables.
@@
Finally, we measured the mean execution time per fused gate and total peak memory usage over the forward and backward paths while performing a task of minimizing a sum of expectation values for the observable $O=IXYZ\ldots IXYZ$ (repeated $5$ times) using the same quantum circuit as above, shown in Fig. 3, with $20$ qubits and a batch size of $100$. The mean execution time per fused gate is shown in Fig. 6(a) and the total peak memory usage over $20$ fused gates is shown in Fig. 6(b). We performed the evaluation of the PyTorch native implementation for this numerical experiment using only GH200 because it requires a huge amount of memory to store the state vectors. Also, we evaluated our proposed method on the RTX 4090 and GH200 due to the insufficient memory capacity on the RTX 5070.
As shown in Fig. 6(a), the implementation of our proposed method performs significantly faster than the PyTorch native implementation. The mean execution time per fused gate for both approaches increases linearly with the number of variational gates to be fused, but that for our proposed method increases more gradually. This trend suggests that reducing the number of accesses to global memory by our proposed method contributes to this acceleration. Additionally, for the result of PyTorch native implementation, the increase in execution time is gradual when an Rz gate is added. As shown in Fig. 6(b), the peak memory usage of the PyTorch native implementation shows a steep linear increase with the number of variational gates to be fused, since all input quantum states for the gates are stored during the forward path for later use in the backward path. For our proposed method, on the other hand, the required memory remains almost constant regardless of the number of variational gates to be fused, and is drastically lower than that of the PyTorch native implementation, since it only needs to store the resulting state of the fused variational gates. In addition, Triton fused (mem save) implementation further reduced the peak memory usage by type converting the state vector from float32 to bfloat16 when temporarily storing it in the forward path. As shown in Fig. 6(c), when $10$ or more variational gates are fused, the mean execution time per fused gate of Triton fused (mem save) is $5\%$ to $10\%$ slower than that of Triton fused on the GH200, and $10\%$ to $20\%$ slower on the RTX 4090. Importantly, we can reduce peak memory usage by $40\%$ at the cost of an approximate $10\%$ increase in computation time. This trade-off is practically important, particularly for a consumer GPU with limited memory capacity and bandwidth.
(a) Execution time per fused gate
@@
Figure 6: (a) Execution time per fused gate and (b) total peak memory usage (c) execution time ratio between Triton Fused and Triton Fused (mem save) when minimizing the sum of expectation values at the final quantum state after applying the fused single-qubit gates shown in Fig. 3 to a $100$ fixed batch of random initial states for one epoch.
@@ 4.2 Throughput for a Hardware Efficient Ansatz
Here, we measure the throughput for a task of minimizing a sum of expectation values for the observable $O=IXYZ\cdots IXYZ$ (repeating $IXYZ$ until the length matches the number of qubits) at the final state, made by applying an HEA (shown in Fig. 7 with $d=1$ ), consisting of Rx, Ry, Rz, and CZ gates, to random initial quantum states. To evaluate the throughput, we set the batch size as large as possible to be stored within the GPU memory, and we estimated the throughput using this batch size and the mean execution time across $1,000$ runs. The results are shown in Fig. 8.
In this numerical experiment, to ensure a sufficient number of gates to be fused, we fused the consecutive single-qubit gates acting on not only the same qubit but also on adjacent qubits, as shown in Fig. 7. This is because fusing gates acting only on the same qubit limits the number of gates that can be fused, potentially making it difficult to hide memory access latency and thereby limiting the acceleration effect. Similarly, we also fused consecutive CZ gates.
As shown in Fig. 8 and Fig. 10, for $12$ qubits and above, our proposed method achieved approximately $20$ times higher throughput compared to the PyTorch native implementation. In particular, on the midrange GPU RTX5070 with limited memory bandwidth, our proposed method achieved approximately $30$ times higher throughput. These results demonstrate that implementing gate fusion in both forward and backward paths yields a significantly higher throughput than the PyTorch native implementation.
For reference, we also plot the results of an existing QML library TorchQuantum [torchquantum2024], with a red dotted line in Fig. 8. The implementation of this library uses a matrix-based approach (e.g., for calculating expectation values), making classical simulation for over $15$ qubits infeasible due to memory constraints. On the other hand, the implementation of our proposed method employs a matrix-free approach, similar to the fast quantum circuit simulator “Qulacs" [suzuki2021qulacs]. Furthermore, by recomputing the intermediate states of fused gates during the backward path, we do not need to store them in global memory. As a result, we achieved a throughput of $3.57$ samples/sec for a $29$-qubit HEA by Triton fused on a single GPU GH200 and $1.62$ samples/sec for a $30$-qubit HEA by Triton fused (mem save) on the same GPU, whereas the Triton fused implementation failed at this scale due to memory constraints. This highlights the practical effectiveness of our memory-saving strategy for large-scale classical simulations.
@@
4.3 Classical simulation of $1,000$ layers of HEA using gradient checkpointing
Finally, as a practical benchmark, we perform a large-scale classical simulation of a deep quantum circuit for QML by combining our proposed method with gradient checkpointing in PyTorch (torch.utils.checkpoint) on the GH200. Specifically, we perform the classical simulation of minimizing the sum of expectation values for the observable $O=IXYZ\cdots IXYZ$ ( consisting of $5$ repetitions of $IXYZ$ ) at the final states, generated by applying a $20$-qubit HEA with $d=1,000$ layers (including $60,000$ parameters), as shown in Fig. 7, to random initial quantum states.
@@ 4.3.1 Preliminary experiment to determine checkpoint block size $b$
Analysis of peak memory usage: As shown in Fig. 9(a), for a small $b$ ($1\leq b\leq 5$), all implementations consume a large amount of memory. This is because a small block size implies a large total number of blocks ($d/b$), making the storage of input state vectors for each block dominant, scaling as $O(d/b)$. As $b$ increases, peak memory usage begins to decrease. Then, the minimum peak memory usage for Triton fused is $1,443$ MB at $b=10$, and that for Triton fused (mem save) is $1,032$ MB at $b=20$. This result demonstrates that we successfully reduced peak memory usage by approximately $30\%$ by temporarily storing the output state vectors of fused variational gates as bfloat16. As $b$ increases further, the memory required to store output state vectors for fused gates within each active block becomes dominant, causing the peak memory usage to increase again, scaling as $O(\lceil\frac{l_{\text{var}}}{\alpha}\rceil\cdot b)$. Interestingly, at $b=100$, the peak memory usage for Triton fused is $5,972$ MB, and that for Triton fused (mem save) is $3,045$ MB, which is almost half the size. This result indicates that when the term $\left(\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil\cdot b\right)$ dominates, the effect of reducing peak memory usage by storing compressed state vectors is consistent with theoretical expectations.
Comparison with theoretical analysis: Here, we compare the peak memory usage of our numerical results with the theoretical analysis. To determine $\alpha$, we first focus on the number of fused gates and stored state vectors. As shown in Fig. 7, for three adjacent qubits, we fused the sequence of Rx, Ry, and Rz gates acting on each of the three adjacent qubits into a single operator ($m=3\times 3=9$). For the remaining two qubits, we fused the corresponding sequence of gates ($m=2\times 3=6$). So, the number of fused variational gates is $\lceil\frac{l_{\text{var}}}{\alpha}\rceil=\lceil\frac{60}{9}\rceil=7$, and each fused variational gate requires storing its output state vector. Note that a fused CZ gate does not require storing its state vectors because it does not have a variational parameter. Since the size of complex64 is $8$ bytes, the theoretical peak memory usage for a $1,000$-layer HEA can be estimated by substituting $n=20$ and $d=1000$ into Eq. (7):
@@
From Fig. 9(a), we can see that the measured values agree well with the theoretical values. In addition, the optimal block size to minimize peak memory usage is $b^{*}=\sqrt{\frac{1000}{7}}\simeq 12$ for the Triton fused implementation, and $b^{*}=\sqrt{\frac{1000}{3.5}}\simeq 17$ for Triton fused (mem save) implementation. These values are close to the measured values ($10$ and $20$, respectively). This consistency means that the peak memory usage behavior of our proposed method is highly predictable, facilitating the estimation of required resources, even for large-scale models.
Execution time and scalability: From Fig. 9(b), the execution time remains almost constant regardless of $b$ within a $1\%$ fluctuation. This result indicates that the cost of recomputation is almost hidden by the memory access latency. These slight fluctuations in the execution time are caused by the overhead of calling the gradient checkpointing function, as well as the recomputation, storage, and type conversion of state vectors within each active block during gradient checkpointing. When $b$ is small (i.e., many checkpoints), it is considered that the overhead of calling the gradient checkpointing function causes a slight increase in execution time. When $b$ is large, the computational cost of recomputing and storing state vectors in each active block grows, causing an increase in execution time. Moreover, for Triton fused (mem save), the type conversion further increases execution time. Notably, the optimal block size $b$ for the fastest execution time for the Triton fused (mem save) implementation is smaller than that for Triton fused. This is because the input state vectors for each checkpoint block during gradient checkpointing are stored without type conversion, whereas the output state vectors of fused variational gates within an active checkpoint block require type conversion, making the optimal block size $b$ for Triton fused (mem save) smaller in terms of total execution time.

### 03 Apr 2026 v2

@@ 4 Numerical experiments
In this section, to evaluate the performance of our proposed method, implemented with Triton [triton2019] and PyTorch [Ansel_PyTorch_2_Faster_2024], we measured the average execution time across multiple runs over a 1,0001{,}000 ms window using the benchmarking utility triton.testing.do_bench in Triton [triton2019]. If a single execution exceeded this duration, the single execution time was evaluated. We also measured peak memory usage using torch.cuda.max_memory_allocated from a single measurement, since peak memory usage remains consistent across multiple executions. Note that although our implementation supports both single precision (float32/complex64) and double precision (float64/complex128), we use float32 (complex64) in our numerical experiments to focus on performance evaluation for practical QML tasks, which often have severe memory constraints. Furthermore, throughout our numerical experiments, we apply quantum circuits to randomly initialized quantum states $|\psi_{0}\rangle$ rather than $|0\rangle^{\otimes n}$ in order to avoid artificial speedups caused by sparsity and to reflect the actual computational load and memory bandwidth utilization.
@@ 4.1 The effect of gate fusion
In this subsection, we investigate the effect of the gate fusion on a quantum circuit consisting of consecutive single-qubit gates, as shown in Fig. 3. Using a $20$-qubit system and a batch size of $100$, we focus on how memory usage and execution time depend on the number of gates to be fused in the forward and backward paths. In the numerical experiments in this subsection, while the PyTorch native implementation sequentially applies $m$ consecutive single-qubit gates acting on each qubit in the repeating order of Rx, Ry, and Rz, our proposed method fuses these $m$ gates into a single operator.
First, we measured the mean execution time to apply these $m$ gates during the forward path. As shown in Fig. 4(a), the execution time of our proposed method increases almost linearly with $m$, but much more gradually than that of the PyTorch native implementation. Consequently, our proposed method achieves a drastic acceleration. For the PyTorch native implementation, we observe that the execution time increases significantly when applying the Rx and Ry gates, whereas it increases gradually only when applying the diagonal matrix Rz gate.
The total peak memory usage after applying the $m$-gate sequence to all $20$ qubits during the forward path is shown in Fig. 4(b). As shown, across all implementations, the peak memory usage remains almost constant regardless of the number of consecutive single-qubit gates to be applied. In the forward path, the peak memory usage of our proposed method (Triton fused) is the lowest among the three implementations. The peak memory usage is 2,5002{,}500 MB, roughly corresponding to three batched state vectors ($3\times 800$ MB). It appears that the input, output, and temporary batched state vectors are held, and the remaining $100$ MB is internal overhead. Similarly, the Triton fused (mem_save) implementation used an additional $400$ MB, caused by storing the batched state vector converted to bfloat16 for the backward path. The PyTorch native implementation consumes approximately $800$ MB more memory than the Triton fused implementation, corresponding to a single batched state vector. This suggests that the PyTorch native implementation temporarily keeps a copy of the batched state vectors during the forward path, likely due to explicitly cloning them for safety. Since PyTorch’s automatic differentiation requires out-of-place operations on inputs to preserve the computation graph, our implementation is considered to be memory-efficient.
@@
(a) Execution time per $m$-gate sequence
@@
Figure 4: (a) Execution time per $m$-gate sequencee and (b) Total peak memory usage when applying $m$ consecutive single-qubit gates on each qubit during the forward path
Next, to evaluate the performance of the backward function itself, we apply the same quantum circuit as above, shown in Fig. 3, and measure the mean execution time per fused backward kernel. From the results shown in Fig. 5(a), for consumer GPUs such as the RTX 5070 and RTX 4090, the computation time of the backward calculation for a fused gate consisting of a few constituent gates is almost constant, i.e., the recomputation time is negligible because the memory access latency to read quantum states and adjoint states from global memory is dominant. As the number of gates to be fused further increases, the computation time begins to scale linearly. On the other hand, for data center GPUs with large memory bandwidth such as the GH200, the computation time increases linearly with the number of gates to be fused from the beginning. Fig. 5(b) shows the total peak memory usage after applying the functions to all $20$ qubits. As shown, the peak memory usage scales linearly but very gradually with the number of gates to be fused, growing by only $30$ KB (from $2516.585$ MB to $2516.613$ MB) even when $m$ is increased to $20$. This indicates that the computation of the fused gates in the backward path is processed efficiently while mitigating the use of temporary variables.
@@
Finally, we measured the mean execution time per $m$-gate sequence and total peak memory usage over the forward and backward paths while performing a task of minimizing a sum of expectation values for the observable $O=IXYZ\ldots IXYZ$ (repeated $5$ times) using the same quantum circuit as above, shown in Fig. 3, with $20$ qubits and a batch size of $100$. The mean execution time per $m$-gate sequence is shown in Fig. 6(a), and the total peak memory usage after applying the consecutive gates to all $20$ qubits is shown in Fig. 6(b). We performed the evaluation of the PyTorch native implementation for this numerical experiment using only GH200 because it requires a huge amount of memory to store the state vectors. Also, we evaluated our proposed method on the RTX 4090 and GH200 due to the insufficient memory capacity on the RTX 5070.
As shown in Fig. 6(a), the implementation of our proposed method performs significantly faster than the PyTorch native implementation. The mean execution time per $m$-gate sequence for both approaches increases linearly with the number of consecutive variational gates, but that for our proposed method increases more gradually. This trend suggests that reducing the number of accesses to global memory by our proposed method contributes to this acceleration. Additionally, for the result of PyTorch native implementation, the increase in execution time is gradual when an Rz gate is added.
As shown in Fig. 6(b), the peak memory usage of the PyTorch native implementation shows a steep linear increase with the number of consecutive gates, since all input quantum states for the gates are stored during the forward path for later use in the backward path. For our proposed method, on the other hand, the required memory remains almost constant regardless of the number of variational gates to be fused, and is drastically lower than that of the PyTorch native implementation, since our proposed method only needs to store the resulting state of the fused variational gates.
In addition, the Triton fused (mem save) implementation further reduced the peak memory usage by type converting the state vector from float32 to bfloat16 when temporarily storing it in the forward path. As shown in Fig. 6(c), when $10$ or more variational gates are fused, the mean execution time per $m$-gate sequence of Triton fused (mem save) is $5\%$ to $10\%$ slower than that of Triton fused implementation on the GH200, and $10\%$ to $20\%$ slower on the RTX 4090. Importantly, we can reduce peak memory usage by $40\%$ at the cost of an approximate $10\%$ increase in computation time. This trade-off is practically important, particularly for a consumer GPU with limited memory capacity and bandwidth.
Note that in Fig. 6(a) and Fig. 6(c), local fluctuations (e.g., at $m=4,9$) are observed on the consumer GPU RTX 4090, whereas the data center GPU GH200 exhibits a smoother trend. These performance fluctuations are typical of consumer GPUs and are caused by background OS interrupts, display rendering tasks, and dynamic voltage and frequency scaling (DVFS), which occasionally compete with GPU resources and memory bandwidth.
(a) Execution time per $m$-gate sequence
@@
Figure 6: (a) Execution time per $m$-gate sequence and (b) total peak memory usage (c) execution time ratio between Triton Fused and Triton Fused (mem save) when minimizing the sum of expectation values at the final quantum state after applying the fused single-qubit gates shown in Fig. 3 to a $100$ fixed batch of random initial states for one epoch. Note that, unlike the smoother trend of the data center GPU GH200, local performance fluctuations (e.g., at $m=4,9$) are observed on the RTX 4090, likely due to background OS tasks and dynamic frequency scaling.
@@ 4.2 Throughput for a Hardware Efficient Ansatz
Here, we measure the throughput for a task of minimizing a sum of expectation values for the observable $O=IXYZ\cdots IXYZ$ (repeating $IXYZ$ until the length matches the number of qubits) at the final state, made by applying an HEA (shown in Fig. 7 with $d=1$ ), consisting of Rx, Ry, Rz, and CZ gates, to random initial quantum states. To evaluate the throughput, we set the batch size as large as possible to be stored within the GPU memory, and we estimated the throughput using this batch size and the mean execution time. The results are shown in Fig. 8.
In this numerical experiment, to ensure a sufficient number of gates to be fused, we fused the consecutive single-qubit gates acting on not only the same qubit but also on adjacent qubits, as shown in Fig. 7. This is because fusing gates acting only on the same qubit limits the number of gates that can be fused, potentially making it difficult to hide memory access latency and thereby limiting the acceleration effect. Similarly, we also fused consecutive CZ gates, as detailed in Appendix B.1.
From the results presented in Fig. 8 and Fig. 10, we observe that for $12$ qubits and above, our proposed method achieved approximately $20$ times higher throughput compared to the PyTorch native implementation. In particular, on the midrange GPU RTX5070 with limited memory bandwidth, our proposed method achieved approximately $30$ times higher throughput. These results demonstrate that implementing gate fusion in both forward and backward paths yields a significantly higher throughput than the PyTorch native implementation.
For reference, we also plot the results of an existing QML library TorchQuantum [torchquantum2024], with a red dotted line in Fig. 8. The implementation of this library uses a matrix-based approach (e.g., for calculating expectation values), making classical simulation for over $15$ qubits infeasible due to memory constraints. On the other hand, the implementation of our proposed method employs a matrix-free approach, similar to the fast quantum circuit simulator “Qulacs" [suzuki2021qulacs]. Furthermore, by recomputing the intermediate states of fused gates within registers during the backward path, we do not need to store them in global memory. As a result, for a $29$-qubit HEA on a single GPU GH200, the Triton fused implementation achieved a throughput of $3.57$ samples/sec, while the Triton fused (mem save) implementation achieved $3.25$ samples/sec. More importantly, the Triton fused (mem save) implementation successfully scaled to a $30$-qubit HEA with a throughput of $1.62$ samples/sec, whereas the Triton fused implementation failed at this scale due to memory constraints. This highlights the practical effectiveness of our memory-saving strategy for large-scale classical simulations.
@@
4.3 Classical simulation of 1,0001{,}000 layers of HEA using gradient checkpointing
Finally, as a practical benchmark, we perform a large-scale classical simulation of a deep quantum circuit for QML by combining our proposed method with gradient checkpointing in PyTorch (torch.utils.checkpoint) on the GH200. Specifically, we perform the classical simulation of minimizing the sum of expectation values for the observable $O=IXYZ\cdots IXYZ$ (consisting of $5$ repetitions of $IXYZ$) at the final states, generated by applying a $20$-qubit HEA with d=1,000d=1{,}000 layers (including 60,00060{,}000 parameters), as shown in Fig. 7, to random initial quantum states.
@@ 4.3.1 Preliminary experiment to determine checkpoint block size $b$
Analysis of peak memory usage: As shown in Fig. 9(a), for a small $b$ ($1\leq b\leq 5$), all implementations consume a large amount of memory. This is because a small block size implies a large total number of blocks ($d/b$), making the storage of input state vectors for each block dominant, scaling as $O(d/b)$. As $b$ increases, peak memory usage begins to decrease. Then, the minimum peak memory usage for Triton fused is 1,4431{,}443 MB at $b=10$, and that for Triton fused (mem save) is 1,0321{,}032 MB at $b=20$. This result demonstrates that we successfully reduced peak memory usage by approximately $30\%$ by temporarily storing the output state vectors of fused variational gates as bfloat16. As $b$ increases further, the memory required to store output state vectors for fused gates within each active block becomes dominant, causing the peak memory usage to increase again, scaling as $O(\lceil\frac{l_{\text{var}}}{\alpha}\rceil\cdot b)$. Interestingly, at $b=100$, the peak memory usage for Triton fused is 5,9725{,}972 MB, and that for Triton fused (mem save) is 3,0453{,}045 MB, which is almost half the size. This result indicates that when the term $\left(\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil\cdot b\right)$ dominates, the effect of reducing peak memory usage by storing compressed state vectors is consistent with theoretical expectations.
Comparison with theoretical analysis: Here, we compare the peak memory usage of our numerical results with the theoretical analysis. To determine $\alpha$ for the Triton fused implementation, we first focus on the number of fused gates and stored state vectors. As shown in Fig. 7, for groups of three adjacent qubits, we fused the sequence of Rx, Ry, and Rz gates acting on each of the three adjacent qubits into a single operator ($\alpha=m=3\times 3=9$). For the remaining two qubits, we fused the corresponding sequence of gates into a smaller operator ($m=2\times 3=6$). So, the number of fused variational gates is $\lceil\frac{l_{\text{var}}}{\alpha}\rceil=\lceil\frac{60}{9}\rceil=7$, and each fused variational gate requires storing its output state vector. Note that a fused CZ gate does not require storing its state vectors because it does not have a variational parameter. Since the size of complex64 is $8$ bytes, the theoretical peak memory usage for a 1,0001{,}000-layer HEA can be estimated by substituting $n=20$ and $d=1000$ into Eq. (7):
@@
From Fig. 9(a), we can see that the measured values agree well with the theoretical values. This consistency indicates that the peak memory usage behavior of our proposed method is highly predictable, facilitating the estimation of required computational resources, even for large-scale models. Specifically, the theoretical optimal block sizes to minimize peak memory usage are $b^{*}=\sqrt{\frac{1000}{7}}\simeq 12$ for the Triton fused implementation and $b^{*}=\sqrt{\frac{1000}{3.5}}\simeq 17$ for the Triton fused (mem save) implementation. These values are close to our empirical observations, where the minimum memory usage were measured at $b=10$ and $b=20$, respectively.
Execution time and scalability: From Fig. 9(b), the execution time remains almost constant regardless of $b$ within a $1\%$ fluctuation. This is fundamentally because the total number of executions of fused gates remains constant regardless of $b$. These slight fluctuations in the execution time are caused by the overhead of calling the gradient checkpointing function, by reading and writing the input quantum states for the gradient checkpoints, and by the cache efficiency. Specifically, when $b$ is small (i.e., many checkpoints), the overhead of calling the gradient checkpointing function, reading and writing input quantum states for the gradient checkpoints, causes a slight increase in execution time. When $b$ is large, an increase in execution time may be attributed to a lower cache efficiency because a larger $b$ requires reading and writing more intermediate quantum states. Moreover, for Triton fused (mem save), the type conversion further increases execution time. Notably, the optimal block size $b$ for the fastest execution time for the Triton fused (mem save) implementation is smaller than that for Triton fused. This is because the input state vectors for each checkpoint block during gradient checkpointing are stored without type conversion, whereas the output state vectors of fused variational gates within an active checkpoint block are stored with type conversion, making the optimal block size $b$ for Triton fused (mem save) smaller in terms of total execution time.